### 15 Dec 2023 v2

# DiffBind: An SE(3) Equivariant Network for Accurate All-Atom Protein-Ligand Docking with Full Pocket Side Chain Flexibility | Jintao Zhu, Zhonghui Gu, Jianfeng Pei, and Luhua Lai |

DiffBind: An SE(3) Equivariant Network for Accurate All-Atom Protein-Ligand Docking with Full Pocket Side Chain Flexibility |
Jintao Zhu, Zhonghui Gu, Jianfeng Pei, and Luhua Lai |
Molecular docking, a key technique in structure-based drug design (SBDD), plays a pivotal role in hit identification for specific targets. Accurate prediction of protein-ligand binding mode is important for precise scoring and rational molecular optimization. Notwithstanding its significance, modelling precise and physically plausible binding conformations is a largely unsolved problem in the real-world docking scenario. Flexible docking is a daunting task as modeling protein conformation changes upon ligand binding is extremely computationally expensive and inaccurate. Currently available deep learning docking methods ignore protein flexibility and fail to ensure the physical plausibility and detailed interactions. In this study, we present DiffBind, a comprehensive full-atom diffusion-based flexible docking model that operates over the product space of ligand movements (translation, rotation, and torsion) and pocket side chain torsion changes. Evaluations reveal that DiffBind has higher accuracy in producing native-like binding structures with physically plausible and detailed interactions than existing docking methods. Furthermore, in the Apo and AlphaFold2 modeled structures, DiffBind demonstrates superior advantages in accurate ligand pose prediction and effective pocket side chain refinement, making it suitable for Apo and AlphaFold2 structure-based drug design. DiffBind should be useful for modeling the pocket-ligand binding structure with significant side chain flexibility. |
1 Introduction
The primary paradigm of drug discovery involves identifying and designing molecules that target key proteins within disease pathways. Historically, screening compound libraries using biochemical platforms has been the predominant approach for identifying novel drugs 1. Since the 1990s, high-throughput screening (HTS) has been employed on libraries ranging from 500,000 to 10 molecules 2, 3, leading to the discovery of several drugs. While the HTS libraries represent a significant advancement over traditional lab-designed ones, they encompass only a fraction of potential drug-like molecules 4. Given the challenges and expenses associated with synthesizing such a vast chemical space, computational methods for screening virtual libraries are frequently employed in drug discovery, allowing exploration of chemical spaces comprising tens of billions of molecules or even more 5, 6, 7.
Structure-based virtual screening (SBVS) enables rapid and cost-effective modeling of target-molecule binding structures from large-scale compound libraries, as well as the evaluation of their binding affinities, thereby identifying potential hits 8, 9, 10. Molecular docking is one of the most frequently employed techniques for SBVS. It is utilized to predict ligand binding poses, characterize protein-ligand binding strength, and identify key interactions 11, 12. In general, conventional docking softwares, including AutoDock4 13, AutoDock Vina 14, 15, Smina 16, Glide 17, and GOLD 18, leverage heuristic search algorithms, to explore a variety of potential ligand conformations. Scoring functions with simplified terms are utilized for the fast estimation of the binding affinity and the priority of ligand poses.
Classical molecular docking methods describe protein-ligand interactions based on the lock-and-key model 19, wherein a rigid receptor binding pocket serves as the "lock" and the molecular docking algorithm primarily optimizes the ligand’s conformation to find a complementary "key". Such rigid receptor docking methods, for the trade-off between accuracy and computational efficiency, always strive to determine the optimal and complementary binding conformation. When known complex structures are available, and ligand molecules are removed and then re-docked into native Holo pockets, rigid receptor docking often achieves impressive success rate 20. However, in real-world cross-docking scenarios, binding pockets may undergo different structural alterations upon various ligands binding 21. Even when ligand-bound structures (Holo) are unavailable for investigated targets, drug screening and design on its unbound state (Apo) or computational modeled structures usually gives unsatisfied hit rate 22, 23, 24, 25, 26, 27. If a docking method does not account for pocket flexibility, its performance can drastically decrease in such cases 28, which may cause potential drugs get ruled out during the early stage of drug discovery. Although AlphaFold2 29 is capable of accurately modeling target protein structures, traditional docking methods that overlook potential side chain flexibility may still fail to perform effectively when applied to these structures30.
Currently, there are two main strategies to address the flexibility of protein pockets. The first involves inducing local conformational rearrangements of the target using force fields or scoring functions. For instance, rDock 31 allows movements of functional groups that result in hydrogen bonds including -OH and -NH3. AutoDockFR (AutoDock for Flexible Receptors) 32 allows users to specify up to 14 flexible side chains in advance and samples reasonable side chain dihedral angles from a rotamer library. Despite its better performance than AutoDock Vina in cross-docking experiments with Apo structures, it is considerably time-consuming and requires prior knowledge of potentially critical side chains in the pocket, which limits its application in SBVS. The second approach is the recently developed deep learning-based methods 33, which coarsen the representation of protein pockets, typically encoding only the protein’s backbone atoms. This representation is not sensitive to minor pocket backbone flexibility and side chain adaptability. Earlier works, like DeepDock 34, TankBind 35, and EDM-Dock 36, predicted pocket residue-molecule distance map, which is used to reconstruct the binding structure. Leveraging powerful equivariant neural networks like EGNN 37, geometric deep learning 38 models such as EquiBind 39, LigPose 40, E3Bind 41, Uni-Mol 42, and KarmaDock 43 could iteratively predict the three-dimensional coordinates of ligands directly around the whole protein (blind docking) or predefined pocket. Recent SOTA blind docking method DiffDock 44, based on the diffusion generative modelling 45, employed the SE(3) equivariant neural network 46 to denoise the rotation, translation, and bond torsion of ligand, and then rank poses by additional confidence model. However, existing deep learning-based docking approaches face limitations in effectively handling protein flexibility. In addition, the generated ligand poses are often implausible for deep learning-based methods 47. Not only are there issues with bond lengths, angles, and torsion angles, leading to high intra energies of the ligand, but also optimizing the conformations generated by RDKit alignment 43, 48 also cannot alleviate clashes between the ligand and protein. Furthermore, ignoring the flexibility and invalidity of ligand poses makes it challenging for these deep learning-based methods to capture key interactions during cross-docking 47. Such short-comings impede subsequent steps, including experts post-optimizing ligands based on these detailed interactions or conducting further studies through molecular dynamics simulations.
Early Apo-Holo pair analysis has shown general consensus that upon ligand binding protein pocket undergoes significant side chain conformation heterogeneity while backbone is relatively rigid in most cases 49, 50, 51. Therefore, in most cases, side chain flexibility modelling is enough for flexible docking. In this study, we developed a full-atom flexible docking model, DiffBind, based on the diffusion framework (Fig.1). We formulated flexible docking as a problem of learning the joint denoising process of four variables in their tangent space: ligand rotation , translation , rotatable bond torsion , and pocket side chain torsion . Following the VE-SDE (variance exploding stochastic differential equation) paradigm 52, starting from the crystal complex , the forward process of the diffusion model, , involves uniformly and continuously sampling time step and injecting noise to the four kinds of movement operator to achieve binding structure perturbation. DiffBind is an SE(3)-equivariant generative model, following the message-passing paradigm 53 of graph neural network, and encodes the intricate interactions between the full-atom pocket and ligand, predicting the scores 52. At the docking procedure, starting from the randomly initialized binding conformation, the scores predicted by DiffBind are used to solve the reverse VE-SDE process 52 to implement denoising sampling. With physics-based scoring function Smina 16 or mixture density neural network (MDN) 34 serving as confidence model, binding structures sampled by DiffBind can be ranked, and then the top-1 complex pose can be selected as the final prediction. In the comprehensive evaluation, starting from pocket conformations with randomized side chain torsion angles, DiffBind outperforms state-of-the-art (SOTA) deep learning techniques and traditional docking methods. DiffBind can not only accurately recover protein pocket side chain conformations, but also generate precise and highly physically plausible ligand binding poses. In cross-docking benchmark, DiffBind, due to its explicit consideration of full pocket side chain flexibility, significantly outperfoms traditional flexible docking methods like AutoDock VinaFlex 32 and rDock 31. It also demonstrated superior performance in generating both accurate and valid binding structures compared to existing docking methods in the Apo and AlphaFold2 modeled structures.
2 Results and discussion
2.1 Performance on the PDBBind time-split test set
The performance of DiffBind is initially assessed using the PDBBind time-split test set 39, 54. We employed four metrics, including ligand Root Mean Square Deviation (L-RMSD), geometric center deviation (C-Dist), proportion of pocket residues with <15 () and side chain Root Mean Square Deviation (sc-RMSD) for the flexible docking evaluation. As is depicted from Fig.2 (DiffBind-best), among the best poses (with lowest L-RMSD) from 40 DiffBind-generated poses for each complex, 76.0 of the ligand poses achieve successful docking (L-RMSD 2 Å). Of these successful DiffBind-best poses, 79.9 correctly identify the binding sites (C-Dist <1 Å); 62.3 demonstrate effective recovery (more than 75 within ); 77.4 exhibit reliable side chain recovery (sc-RMSD <1 Å). The results of ablation experiments conducted on the hyperparameters related to network sampling and denoising within the DiffBind framework are detailed in Supplemental Fig. S3. However, DiffBind-best represents the optimal scenario achievable by the DiffBind model, assuming a perfect confidence model could identify all the best poses. To enhance pose selection, we have developed two confidence models for ranking the generated poses (Fig.3.(A)). The first model (DiffBind-Smina) employs the all-atom physics-based scoring function Smina 16, and the second one (DiffBind-MDN) utilizes a MDN (Mixture Density Network) network (Fig. S2) trained on the PDBbind time-split training set by this work. Using Smina to select the top-1 binding poses, DiffBind-Smina attains a success rate of 51.2 in the PDBbind time-split test set. Among the top-1 poses ranked by Smina, 59.8 can accurately locate binding positions; 58.7 show effective recovery; 74.7 have reliable side chain recovery. The results demonstrate that DiffBind can accurately reconstruct the side chain conformations consistent with experimental pocket-ligand interactions, allowing the Smina to effectively select high-quality binding poses. When ranking sampled poses via the MDN model, DiffBind-MDN achieves a little bit lower success rate of 46.3 compared to DiffBind-Smina. Among the top-1 poses ranked by MDN confidence model, 57.3 correctly identify binding positions; 62.3 demonstrate effective recovery; 75.2 have reliable side chain recovery.
Subsequently, the performance of traditional and recent deep learning-based methods is evaluated for comparison. The top-1 docking poses of each method, selected based on its confidence model or scoring function, are analyzed. DiffBind-Smina and DiffBind-MDN significantly outperform other deep learning-based pocket docking methods (Fig. 3.(B)), including KarmaDock with RDKit 48 ligand conformation alignment (KarmaDock Align) and TankBind with predefined pocket (TankBind-Pocket). This highlights the advantage of our full-atom based model. Even when compared to the traditional rigid receptor docking methods AutoDock Vina and Glide, with the experimentally determined side chain conformations (redock), DiffBind-Smina achieves a marginally higher success rate without knowing the side chain conformations in the complex. We further used Rosetta 55 to repack side chain conformations in these Holo structures to simulate an Apo-like state for each target protein. The docking success rate of Vina and Glide significantly decreases in these Apo-like proteins, underscoring the limitations of rigid receptor docking methods in handling side chain movements and the importance of flexible docking in virtual screening. To illustrate the challenges of flexible docking, we re-trained the KarmaDock by integrating a ResNet module 56 (Supplemental Fig. S4) for predicting side chain torsion angles, resulting in a new model named KarmaDock-sc Align. DiffBind significantly surpasses KarmaDock-sc Align in terms of side chain recovery (Fig. 3.(C)). Compared to KarmaDock, the performance of KarmaDock-sc Align significantly declines (Fig. 3.(B)) due to the difficulty in balancing ligand coordinate recovery with side chain torsion recovery, highlighting the complexity inherent in flexible docking. As is widely recognized, factors like the number of heavy atoms and rotatable bonds in a ligand profoundly impact the success rate of conventional docking programs 57. Hence, we examine the relationship between L-RMSD from the DiffBind model and various ligand characteristics, such as molecular weight, rotatable torsion bonds, and changes in solvent accessible surface area (DeltaSASA) upon binding. Contrary to traditional methods 17, 14, 31, 58, as the ligand complexity and size increase, the performance of DiffBind-Smina does not deteriorate significantly (Fig.3.(D)).
2.2 Performance on the Posebusters test set
To evaluate the physical plausibility of poses generated by DiffBind, we compared its performance to other baseline methods using the Posebusters test set and the Posebusters suite 47, a tool designed to assess the validity of ligand-protein complexes based on criteria including bond length, planarity of aromatic rings in ligands, and clashes between ligands and proteins. Success in docking is redefined as a pose having an L-RMSD less than 2 Å and simultaneously passing the physical validity check by Posebusters, with this success rate termed as the PB-success rate. The Posebusters test set comprises 428 distinct complexes released since 2021, with no overlap with the PDBBind v2020 dataset. To demonstrate that the success rate of DiffBind is not solely due to local ligand energy relaxation and its superior in side chain packing, the performance of other methods is evaluated with stricter energy minimization for the ligand, given the experimental side chain conformation. For poses generated by DiffBind, ligand energy minimization is conducted using the side chains as predicted by the model. The energy minimization is performed using the AMBER ff14sb force field 59 for prorein and the Sage force field 60 for ligand in OpenMM 61, as used in the Posebusters paper 47. Fig.4.(A) shows that traditional rigid receptor docking methods like Glide perform best on re-docking when provided with the correct Holo pocket environment, followed by Vina and Gold, with most of their generated docking poses being physically valid. However, their performance significantly deteriorates when docking with Rosetta-repacked proteins, highlighting their heavy dependence on side chain conformations. Traditional flexible docking methods rDock and VinaFlex are also involved in comparison. VinaFlex, heavily reliant on predefined flexible side chains, performs the worst in our scenario where information about flexible side chains is assumed unavailable. rDock, capable of optimizing functional groups prone to forming hydrogen bonds in side chains, achieves higher success rate in repacked proteins compared to Vina and Glide, but lower success rate in proteins with ground-truth side chains. For these traditional methods, their PB-success rate is only slightly lower than their overall success rate, indicating that most generated poses is validated by Posebusters suite due to the physical components in their scoring functions. Therefore, the post ligand optimization using force field does not cause obvious impact to their PB-success rate. Among blind docking methods, DiffDock shows better performance (success rate of 38) than TankBind (16) and EquiBind (2), but most of their generated poses are invalid due to ignoring protein side chains (Fig.4.(A)). Ligand energy minimization significantly improves the PB-success rate of DiffDock (35). TankBind and EquiBind also see improvements in PB-success rate with energy minimization, but still lag behind DiffDock (Fig.4.(B)). Although blind docking is a tough task for its broad searching space in the whole protein, but flexible docking like DiffBind, denoising a chaotic side chain conformation into a well-packed conformation having valid interaction with the ligand, has much more objectives for prediction. DiffBind, utilizing Smina scoring function or MDN network as the confidence model to select the top-1 pose from 40 generated ones, outperforms all other deep learning-based methods in both blind and pocket docking. DiffBind-Smina and DiffBind-MDN demonstrate both high success rate (50.2 for DiffBind-Smina and 48.1 for DiffBind-MDN) and PB-success rate (49.1 for DiffBind-Smina and 44.4 for DiffBind-MDN), with lower penalties by Posebusters compared to other deep learning-based methods, showcasing the capability of DiffBind in generating accurate and physically plausible complex poses. The performance of DiffBind is comparable to traditional rigid receptor docking methods using known ground-truth side chain conformations for redocking. As is depicted from Supplemental Fig. S6, DiffBind shows its effectiveness in binding site identification and pocket side chain recovery on Posebusters test set, as well. Force field optimization has minimal impact on DiffBind generated structures, which also demonstrates the high physical plausibility of DiffBind generated poses. Among other deep learning-based pocket docking methods, KarmaDock Align achieves the highest success rate (30.4) but a very low PB-success rate (6.1). Force field optimization of ligands rescues most poses with L-RMSD 2 Å into good physical validity. KarmaDock Align, EDM-Dock, TankBind-pocket, Uni-Mol, and EDM-Dock, which focus on fitting the RMSD of the ligand during training, tend to ignore the intra energy of the generated poses and protein side chains, as is shown from supplemental Fig. S8. Indeed, force field optimization is not allowed in realistic docking to meet the demands of high-throughput screening.
Four specific cases from the Posebusters test set (supplementary Fig. S9), never trained or seen by DiffBind, are presented to highlight its superiority over other methods focusing solely on ligand coordinates while neglecting ligand conformation validity. In complexes with PDB ID 6TW5, 7PRM, 7T1D, and 7CD9, DiffBind successfully docks ligands into precise positions with valid conformations and recovers pocket side chains into good interaction with ligands. In contrast, KarmaDock Align, EDM-Dock, and TankBind-pocket fails to predict correct binding ligand poses, and their generated poses cannot pass the physical plausibility check of the Posebusters suite. As is shown from Supplementary Table. S4, ligand poses generated by EDM-Dock and TankBind-pocket exhibit both internal invalidity (including internal steric clash, bump aromatic ring, etc) and steric clash with proteins, while KarmaDock Align, due to using RDKit for ligand pose alignment, frequently fails in ligand-protein clash.
| Method | Ensemble-CDK2 | Ensemble-EGFR | Ensemble-FXA | ApoRef | CASF2016 | GPCR-AF2 | |||||||||||||
RMSD Mean / Å |
RMSD Median / Å |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
Traditional rigid receptor docking methods |
Vina |
6.12±2.63 |
6.26 |
0.079 |
6.37±2.34 |
6.90 |
0.060 |
4.35±3.17 |
3.02 |
0.344 |
6.98±3.35 |
7.22 |
0.082 |
4.93±3.62 |
4.46 |
0.294 |
7.42±3.99 |
7.77 |
0.136 |
| LinF9 |
6.12±2.50 |
6.31 |
0.061 |
6.31±2.43 |
6.63 |
0.119 |
3.88±2.90 |
2.96 |
0.365 |
6.59±3.08 |
6.67 |
0.089 |
4.99±3.56 |
4.52 |
0.272 |
5.66±2.54 |
5.18 |
0.076 |
| Smina |
6.15±2.65 |
6.28 |
0.079 |
6.33±2.53 |
6.90 |
0.090 |
4.17±3.16 |
2.70 |
0.362 |
7.08±3.47 |
7.43 |
0.077 |
4.88±3.69 |
4.38 |
0.304 |
7.25±4.15 |
7.41 |
0.152 |
| Gnina |
5.67±2.63 |
5.78 |
0.099 |
6.02±2.69 |
6.29 |
0.090 |
3.83±2.99 |
2.44 |
0.394 |
6.67±3.65 |
6.51 |
0.082 |
4.57±3.56 |
3.74 |
0.320 |
6.72±3.95 |
6.06 |
0.106 |
| Glide |
4.81±2.60 |
5.21 |
0.154 |
6.87±3.63 |
7.67 |
0.090 |
3.49±2.72 |
2.31 |
0.271 |
6.62±3.51 |
6.85 |
0.091 |
4.28±3.20 |
3.74 |
0.219 |
5.37±3.63 |
4.57 |
0.182 |
Traditional flexible docking methods |
VinaFlex |
8.04±2.92 |
7.88 |
0.013 |
9.03±2.47 |
9.446 |
0.000 |
9.19±2.22 |
9.42 |
0.005 |
8.81±2.94 |
9.06 |
0.015 |
8.80±2.98 |
9.07 |
0.023 |
7.36±3.33 |
7.07 |
0.045 |
| rDock |
4.61±2.76 |
4.62 |
0.257 |
5.45±2.98 |
4.70 |
0.134 |
3.29±2.66 |
2.28 |
0.440 |
5.46±3.02 |
5.31 |
0.157 |
4.11±2.82 |
3.58 |
0.296 |
5.37±3.60 |
5.12 |
0.212 |
Deep learning-based docking methods |
TankBind-pocket |
2.17±1.85 |
1.62 |
0.100 |
2.21±0.96 |
1.82 |
0.015 |
1.73±0.96 |
1.50 |
0.110 |
2.63±1.88 |
1.97 |
0.040 |
2.37±2.38 |
1.63 |
0.123 |
3.74±1.98 |
3.42 |
0.015 |
| EDM-Dock |
2.62±1.27 |
2.32 |
0.051 |
2.95±1.05 |
2.66 |
0.015 |
2.68±0.98 |
0.45 |
0.009 |
3.33±1.55 |
3.09 |
0.011 |
3.06±1.92 |
2.55 |
0.064 |
4.72±1.95 |
4.61 |
0.00 |
| KarmaDock Align |
1.89±1.20 |
1.58 |
0.135 |
2.65±2.06 |
1.93 |
0.045 |
2.43±0.92 |
2.30 |
0.009 |
2.47±1.56 |
2.04 |
0.047 |
2.36±1.57 |
1.89 |
0.136 |
4.30±2.20 |
3.91 |
0.015 |
| DiffBind-Smina |
2.31±1.80 |
1.73 |
0.564 |
3.44±2.53 |
2.57 |
0.403 |
1.54±1.14 |
1.30 |
0.789 |
2.96±2.32 |
2.10 |
0.434 |
2.30±2.28 |
1.58 |
0.566 |
4.92±4.12 |
4.02 |
0.227 | |
| DiffBind-MDN |
1.85±1.34 |
1.48 | 0.674 |
2.58±2.15 |
1.80 | 0.478 |
1.42±0.70 |
1.35 |
0.794 |
2.32±1.73 |
1.76 | 0.476 |
1.87±1.93 |
1.25 | 0.636 |
3.64±2.59 |
2.74 |
0.212 |
Best performance in bold. RMSD Mean and RMSD Medium, lowest; PB-SR, highest. RMSD Mean and RMSD Median denote the average ± standard deviation and median of L-RMSD for top-1 generated ligand poses from each complex, respectively. PB-SR denotes PB-success rate.