### 19 Dec 2023 v3

# DiffBindFR: An SE(3) Equivariant Network for Flexible Protein-Ligand Docking | Jintao Zhu, Zhonghui Gu, Jianfeng Pei, and Luhua Lai |

DiffBindFR: An SE(3) Equivariant Network for Flexible Protein-Ligand Docking |
Jintao Zhu, Zhonghui Gu, Jianfeng Pei, and Luhua Lai |
Molecular docking, a key technique in structure-based drug design, plays pivotal roles in protein-ligand interaction modeling, hit identification and optimization, in which accurate prediction of protein-ligand binding mode is essential. Conventional docking approaches perform well in redocking tasks with known protein binding pocket conformation in the complex state. However, in real-world docking scenario without knowing the protein binding conformation for a new ligand, accurately modeling the binding complex structure remains challenging as flexible docking is computationally expensive and inaccurate. Typical deep learning-based docking methods do not explicitly consider protein side chain conformations and fail to ensure the physical plausibility and detailed atomic interactions. In this study, we present DiffBindFR, a full-atom diffusion-based flexible docking model that operates over the product space of ligand overall movements and flexibility and pocket side chain torsion changes. We show that DiffBindFR has higher accuracy in producing native-like binding structures with physically plausible and detailed interactions than available docking methods. Furthermore, in the Apo and AlphaFold2 modeled structures, DiffBindFR demonstrates superior advantages in accurate ligand binding pose and protein binding conformation prediction, making it suitable for Apo and AlphaFold2 structure-based drug design. DiffBindFR provides a powerful flexible docking tool for modeling accurate protein-ligand binding structures. |
1 Introduction
The primary paradigm of drug discovery involves identifying and designing molecules that target key proteins within disease pathways. Historically, screening compound libraries using biochemical platforms has been the predominant approach for identifying novel drugs 1. Since the 1990s, high-throughput screening (HTS) has been employed on libraries ranging from 500,000 to 10 molecules 2, 3, leading to the discovery of several drugs. While the HTS libraries represent a significant advancement over traditional lab-designed ones, they encompass only a fraction of potential drug-like molecules 4. Given the challenges and expenses associated with synthesizing such a vast chemical space, computational methods for screening virtual libraries are frequently employed in drug discovery, allowing exploration of chemical spaces comprising tens of billions of molecules or even more 5, 6, 7.
Structure-based virtual screening (SBVS) enables rapid and cost-effective modeling of target-molecule binding structures from large-scale compound libraries together with the evaluation of their binding affinities for identifying potential hits 8, 9, 10. Molecular docking is one of the most frequently employed techniques for SBVS, which is utilized to predict ligand binding poses, characterize protein-ligand binding strength, and identify key interactions 11, 12. In general, conventional docking approaches, including AutoDock4 13, AutoDock Vina 14, 15, Smina 16, Glide 17, and GOLD 18, leverage heuristic search algorithms, to explore a variety of potential ligand conformations. Scoring functions with simplified terms are utilized for fast estimation of binding affinity and priority of ligand poses.
Classical molecular docking methods describe protein-ligand interactions based on the lock-and-key model 19, wherein a rigid receptor binding pocket serves as the "lock" and the molecular docking algorithm primarily optimizes the ligand’s conformation to find a complementary "key". Such rigid receptor docking methods, for the trade-off between accuracy and computational efficiency, strive to determine the optimal and complementary binding conformation. When known complex structures are available, and ligand molecules are removed and then re-docked into the native Holo pockets, rigid receptor docking often achieves impressive success rate 20. However, in real-world docking tasks without knowing the binding conformation in advance, ligand induced pocket conformational changes may produce wrong docking results 21. Virtual screening and rational design on the unbound (Apo) or computationally modeled structures usually give unsatisfied hit rate 22, 23, 24, 25, 26, 27. Without accounting for pocket flexibility, the performance of docking methods experiences drastic decrease in such cases 28, which may rule out potential hits during the early stage of drug discovery. Although AlphaFold2 29 is capable of accurately modeling target protein structures, traditional docking methods that overlook potential side chain flexibility perform less effectively when applied to the predicted structures30.
Currently, there are two main strategies to address the flexibility of protein pockets. The first approach involves inducing local conformational rearrangements of the target using force field or scoring function-based calculations. For instance, rDock 31 allows movements of functional groups that can form hydrogen bonds including -OH and -NH3. AutoDockFR (AutoDock for Flexible Receptors) 32 allows users to specify up to 14 flexible side chains in advance and samples reasonable side chain dihedral angles from a rotamer library. Despite its better performance than AutoDock Vina in cross-docking benchmarks with Apo structures, it is considerably time-consuming and requires prior knowledge of potentially critical side chains in the pocket, which limits its application in SBVS. The second approach is the recently developed deep learning-based methods 33, which coarsen the representation of protein pockets by only encoding the protein backbone atoms without explicitly including side chain atoms. This kind of representation is insensitive to minor pocket backbone flexibility and side chain adaptability. Earlier works, like DeepDock 34, TankBind 35, and EDM-Dock 36, predicted pocket residue-ligand distance map, which is used to reconstruct the binding structure. Leveraging powerful equivariant neural networks like EGNN 37, geometric deep learning 38 models such as EquiBind 39, LigPose 40, E3Bind 41, Uni-Mol 42, and KarmaDock 43 iteratively predict the three-dimensional coordinates of ligands directly around the whole protein (blind docking) or predefined pocket. Recent SOTA blind docking method DiffDock 44, based on the diffusion generative modelling 45, employed the SE(3) equivariant neural network 46 to denoise the rotation, translation, and bond torsion of ligand, and then rank poses by additional confidence model. However, these existing deep learning-based docking approaches face limitations in effectively handling protein flexibility and the generated ligand poses are often implausible 47. The generated ligand structures often contain clashes with the target and irrational bond lengths, angles, and torsion angles that lead to high intra energies. Ligand conformational optimization using tools such as RDKit alignment 43, 48 cannot completely alleviate ligand and protein clashes. Furthermore, ignoring the target flexibility and validity of ligand poses makes it challenging for these deep learning-based methods to capture key interactions in docking 47. Recently, building upon the methodology of optimizing ligand coordinate recycling as developed in LigPose and KarmaDock, a deep learning-based flexible docking method named FlexPose 49 has extened its predictive capabilities of pocket side chain coordinates. This advancement allows for more details of interaction information in cross-docking applications. However, like LigPose and KarmaDock, FlexPose encounters the same inherent limitation. The methodological focus on fitting coordinates within Euclidean space tends to overfit the overall RMSD (Root Mean Square Deviation). Consequently, FlexPose, akin to its predecessors, is inevitably limited by conformational rationality. Overall, these inadequacies of current methods impede subsequent steps, such as post-optimization of ligands by experts based on the detailed interactions or conducting further studies through molecular dynamics simulations.
Early Apo-Holo pair analysis has shown general consensus that upon ligand binding protein pocket undergoes significant side chain conformation heterogeneity while backbone is relatively rigid in most cases 50, 51, 52. Therefore, in most cases, side chain flexibility modelling is enough for flexible docking. In this study, we developed a full-atom flexible docking model, DiffBindFR, based on the diffusion framework (Fig.1). In the comprehensive evaluation, starting from pocket conformations with randomized side chain torsion angles, DiffBindFR outperforms state-of-the-art (SOTA) deep learning techniques and traditional docking methods. Owing to the explicit full-atom modeling of pocket residues, and its learning of joint optimization of variables within the entire system across a product space composed of torsional angles, rotations, and translations, DiffBindFR can not only accurately recover protein pocket side chain conformations, but also generate precise and highly physically plausible ligand binding poses. In cross-docking benchmark, DiffBindFR, due to its capability of full pocket side chain optimization, significantly outperforms traditional flexible docking methods like AutoDock VinaFlex 32 and rDock 31. It also gives superior performance in generating both accurate and valid binding structures compared to existing docking methods in the Apo dataset and AlphaFold2 modeled structures.
2 Overview of DiffBindFR
We formulated flexible docking as a problem of learning the joint denoising process of four variables in their tangent space: ligand rotation , translation , rotatable bond torsion , and pocket side chain torsion . Following the VE-SDE (variance exploding stochastic differential equation) paradigm 53, starting from the crystal complex , the forward process of the diffusion model, , involves uniformly and continuously sampling time step and injecting noise to the four kinds of movement operator to achieve binding structure perturbation. DiffBindFR is an SE(3)-equivariant generative model, following the message-passing paradigm 54 of graph neural network, that encodes the intricate interactions between the full-atom pocket and ligand, and predicts the scores 53. In the docking procedure, starting from the randomly initialized binding conformation, the scores predicted by DiffBindFR are used to solve the reverse VE-SDE process 53 to implement denoising sampling. With physics-based scoring function Smina 16 or mixture density neural network (MDN) 34 serving as confidence model, binding structures sampled by DiffBindFR can be ranked, and then the top-1 complex pose can be selected as the final prediction.
2.1 Diffusion generative model
The diffusion model utilize the framework of stochastic differential equations 55 to diffuse the data distribution described as follows:
For , denotes a vector-valued function called the drift coefficient of , and denotes a scalar function called the diffusion coefficient of . The lack of canonical local coordinate system defined for ligand molecules, makes the drift coefficient hard to design for the ligand rotation. Consequently, the drift coefficient is set to be , and the model becomes the score-based generative model 53. The reverse diffusion running backwards in time, which is also known as the denoising process, is given by the following reverse-time SDE:
To estimate , we can train a score-based neural network to fit it. The standard score-match loss function is as follows:
is the precomputed weight factors.
2.2 Pose transformations and diffusion on the product space
We choose the specific SDE for forward diffusion process as follows:
denotes the noises that injected into ligand Rotation , translation , rotatable bond torsion and pocket side chain torsion . According to the specific group that each variable lies in, we would design the form of corresponding carefully for diffusion kernel and the score computation. For a ligand pose with atoms, , translation of a ligand pose lies in the 3D translation group . The diffusion kernel for ligand is a Gaussian function with variance as follows, which is also utilized for computing the score of ligand translation :
As rotation of a ligand pose lies in the 3D rotation group , distribution 56, 57 was chosen as the diffusion kernel. In specific, rotation matrix can be split into a unit vector uniformly as the rotation axis and a axis-angle . Consequently, the functionality of can be replaced by and . The diffusion kernel for ligand rotation is as follows:
The score of rotation diffusion can be computed according to
where is a scalar variance for parameterizing the distribution. Torsion of pocket side chains and ligand rotatable bonds lie in the group and group respectively, where and denote the number of all from the pocket side chain and all from the ligand. Since each torsion angle coordinate lies in , the m torsion angles of a conformer define a hypertorus . We introduced the diffusion kernel from the work of Torsional Diffusion 58 to satisfy angle periodicity, and compute its score as follows:
Torsion of ligand rotatable bonds are dealt with the same way as pocket side chains.
Following the equation(3), the loss function is set as follows:
The forward diffusion and reverse diffusion are both performed in the product space 59 of , corresponding to the aforementioned four kinds of transformation.
During the forward diffusion process, we would sample for each pocket-ligand pair, and then utilize the defined diffusion kernel to sample each transformation. The torsions of ligand and pocket side chains are first applied to the pose, followed by translation and rotation.
The starting point of the denoising stage is a ligand conformation generated by RDKit 48 and pocket side chains, with each type of transformation sampling from their . According to equation(2), we update complex pose using the predicted score for each type of transformation. After applying the translation and rotation matrix constructed from predicted score, torsion angles get updated. It is noteworthy that there exists entanglement between ligand translation/rotation and its bond torsion, ligand pose need to be re-aligned to the pose before bond torsion, which will lead to model-unaware structural alignment error. With the sampled pocket side chains fixed, we perform fast local energy relaxation on the ligand by Smina 16 for error correction, obtaining the final binding conformations. The number of the denoising steps is defined as 22, and 40 poses are sampled for each pocket-ligand pair, which takes in average 50 s when the batch size is set to 16 on a single 32 GB NVIDIA Tesla V100-SXM2 GPU card.
2.3 Confidence model
We have explored two approaches to rank the poses generated by DiffBindFR. First, the traditional scoring function Smina is utilized to quickly score the generated full-atom pocket-ligand poses. Second, a deep learning-based scoring model based on mixture density network (MDN) is trained to fit the distance distribution between ligand atoms and pocket residues. The architecture of our MDN model is similar to the scoring module of KarmaDock, and it shares the similar input representations with DiffBindFR. To better cater for the full-atom complex system, we set the distance pairs as each ligand atom with their nearest atoms from each pocket residues. More details of MDN model can be found in Supplementary Section 3.
3 Datasets and evaluation metrics
3.1 Datasets
3.1.1 PDBbind time-split dataset
We use the PDBbind v2020 dataset 60 for training and evaluation. For each target protein-ligand pair within the PDBbind v2020 dataset, we define the protein pocket as any residues within 12 Å buffer of any heavy atoms in the ligand molecule. Following the time split strategy proposed by the work of EquiBind 39, where 363 complex structures uploaded later than 2019 serve as test set. After removing ligands that exist in the test set, the remaining 16739 structures are used for training and 968 structures are used for validation. The dataset is named as "PDBbind time-split dataset" in the article.
The time-split of PDBbind is supported to be more strict and reasonable with the protein sequence similarity of 0.484 between test set and training&validation set, compared to CASF2016-split whose protein sequence similarity is 1.00 (Supplemental Table. S2). Volkov et al. 61 have shown that the time-split of PDBbind is more practical and critical over artificial splits such as ligand scaffolds or protein sequence/structure similarity for the model generalization in drug repurposing, lead optimization, and virtual screening.
3.1.2 Posebusters test set
The PoseBusters test set 47 is a meticulously curated collection of crystal complexes sourced from the PDB 62. This set encompasses a diverse array of high-caliber, recent protein-ligand complexes characterized by drug-like molecules. With 428 distinct complexes, inclusive of unique proteins and ligands released since 2021, it ensures no overlap with the complexes found in the PDBbind v2020 dataset.
3.1.3 CD test set
Given the current absence of a large-scale benchmark dataset for cross-docking, especially to address various cross-docking scenarios (including Apo-Holo and cross-docking between different Holo states), we have established a benchmark dataset tailored for the cross-docking evaluation, termed CD test set. We integrated ApoRef 24, a test set constructed by constrained MD for inducing Apo-like pockets into Holo-like pockets; several prominent ensemble docking targets including CDK2, EGFR, FXA; CASF2016 60 and GPCR-AF2 30 that contains 18 human GPCR complexes published after April 30, 2018. ApoBind database 63 and AHoJ 64 are utilized to search for corresponding Apo states based on queried protein-ligand pairs, thereby creating pairs for the Apo-Holo and Holo-Holo mixed cross-dock dataset. The detailed protocol for constructing the CD test set can be found in Supplementary Section 1. The finalized CD test set comprises of 14,194 structural pairs for cross-docking benchmark tests.
3.2 Evaluation metrics
We utilize the Ligand Root Mean Square Deviation (L-RMSD) to assess the predictive quality of ligand conformations. Meanwhile, the evaluation of side chain conformations’ predictive quality is based on the side chain Root Mean Square Deviation (sc-RMSD). Let represents the generated ligand pose, and denotes the native ligand pose.
3.2.1 L-RMSD
We take into account Ligand Root Mean Square Deviation (L-RMSD) corrected for symmetry. The precise calculation formula is given below. Herein, represents the number of heavy atoms in the ligand, and denotes the isomers of the ligand molecular graph.
3.2.2 Success rate
L-RMSD 2 Å is widely recognized as a benchmark indication of successful docking 65. In fact, for cross-docking evaluations, the threshold for determining docking success can be relaxed to 2.5 Å. Nonetheless, to ensure equitable comparison, we adhere to the stricter threshold in this context.
3.2.3 PB-success rate
The PoseBusters test suite serves as a rigorous validation tool, assessing both the chemical and geometric consistency of a ligand, inclusive of its stereochemistry. Moreover, it evaluates the physical plausibility of intra-molecular and intermolecular measurements, focusing on factors like the planarity of aromatic rings, canonical bond lengths, and potential protein-ligand clashes. Therefore, the PoseBusters suite provides users a more accurate and realistic estimation of the success rate, PB-success rate, through further checking the physical plausibility of poses with L-RMSD 2 Å.
3.2.4 sc-RMSD
Given that the pocket backbone remains fixed, we compute the RMSD for the side chains of each residue individually and subsequently average the results. Furthermore, in consideration of the symmetrical topology inherent in side chain structures, symmetry corrections have been implemented for the ASP, GLU, PHE, and TYP residues. We regard an sc-RMSD value of less than 1 Å as indicative of success.
4 Results and discussion
4.1 Performance on the PDBbind time-split test set
The performance of DiffBindFR is first assessed on the PDBbind time-split test set 39, 60. We employed two metrics, including ligand Root Mean Square Deviation (L-RMSD) and side chain Root Mean Square Deviation (sc-RMSD) for the flexible docking evaluation. As is depicted from Fig.2.(B) (DiffBindFR-best), among the best poses (with lowest L-RMSD) from 40 DiffBindFR-generated poses for each complex, 76.0 of the ligand poses achieve successful docking (L-RMSD 2 Å). Of these successful DiffBindFR-best poses, 77.4 exhibit reliable side chain recovery (sc-RMSD <1 Å). The results of ablation experiments conducted on the hyperparameters related to network sampling and denoising within the DiffBindFR framework are detailed in Supplemental Fig. S3. DiffBindFR-best represents the optimal scenario achievable by the DiffBindFR model, assuming a perfect confidence model could identify all the best poses. To enhance pose selection, we have developed two confidence models for ranking the generated poses (Fig.2.(A)). The first model (DiffBindFR-Smina) employs the all-atom physics-based scoring function Smina 16, and the second one (DiffBindFR-MDN) utilizes a MDN (Mixture Density Network) network (Fig. S2) trained on the PDBbind time-split training set by this work. Using Smina to select the top-1 binding poses, DiffBindFR-Smina attains a success rate of 51.2 in the PDBbind time-split test set. Among the top-1 poses ranked by Smina, 74.7 have reliable side chain recovery. The results demonstrate that DiffBindFR can accurately reconstruct the side chain conformations consistent with experimental pocket-ligand interactions, allowing the Smina to effectively select high-quality binding poses. When ranking sampled poses via the MDN model, DiffBindFR-MDN achieves a little bit lower success rate of 46.3 compared to DiffBindFR-Smina. Among the top-1 poses ranked by MDN confidence model, 75.2 have reliable side chain recovery.
Subsequently, the performance of traditional and recent deep learning-based methods is evaluated for comparison. The top-1 docking poses of each method, selected based on its confidence model or scoring function, are analyzed. DiffBindFR-Smina and DiffBindFR-MDN significantly outperform other deep learning-based pocket docking methods (Fig. 3.(A)), including KarmaDock with RDKit 48 ligand conformation alignment (KarmaDock Align) and TankBind with predefined pocket (TankBind-Pocket). This highlights the advantage of our full-atom based model. Even when compared to the traditional rigid receptor docking methods AutoDock Vina and Glide, with the experimentally determined side chain conformations (redock), DiffBindFR-Smina achieves a marginally higher success rate without knowing the side chain conformations in the complex. We further used Rosetta 66 to repack side chain conformations in these Holo structures to simulate an Apo-like state for each target protein. The docking success rate of Vina and Glide significantly decreases in these Apo-like proteins, underscoring the limitations of rigid receptor docking methods in handling side chain movements and the importance of flexible docking in virtual screening. To illustrate the challenges of flexible docking, we re-trained the KarmaDock by integrating a ResNet module 67 (Supplemental Fig. S4) for predicting side chain torsion angles, resulting in a new model named KarmaDock-sc Align. DiffBindFR significantly surpasses KarmaDock-sc Align in terms of side chain recovery (Fig. 3.(B)). Compared to KarmaDock, the performance of KarmaDock-sc Align significantly declines (Fig. 3.(A)) due to the difficulty in balancing ligand coordinate recovery with side chain torsion recovery, highlighting the complexity inherent in flexible docking. As is widely recognized, factors like the number of heavy atoms and rotatable bonds in a ligand profoundly impact the success rate of conventional docking programs 68. Hence, we examine the relationship between L-RMSD from the DiffBindFR model and various ligand characteristics, such as molecular weight, rotatable torsion bonds, and changes in solvent accessible surface area (DeltaSASA) upon binding. Contrary to traditional methods 17, 14, 31, 40, as the ligand complexity and size increase, the performance of DiffBindFR-Smina does not deteriorate significantly (Fig.3.(C)).
4.2 Performance on the Posebusters test set
Given that the similarity between samples to be predicted and those used in training can influence the performance of deep learning methods, Buttenschoen et al. 47, aiming for a more equitable comparison with traditional docking methods based on scoring functions, have curated a dataset called Posebusters test set from the PDB database. The Posebusters test set exclusively comprises 428 complexes on which the deep learning methods have not been trained. To evaluate the physical plausibility of poses generated by DiffBindFR, we compared its performance to other baseline methods using the Posebusters test set and the Posebusters suite 47, a tool designed to assess the validity of ligand-protein complexes based on criteria including bond length, planarity of aromatic rings in ligands, and clashes between ligands and proteins. Success in docking is redefined as a pose having an L-RMSD less than 2 Å and simultaneously passing the physical validity check by Posebusters, with this success rate termed as the PB-success rate. The Posebusters test set comprises 428 distinct complexes released since 2021, with no overlap with the PDBbind v2020 dataset. To demonstrate that the success rate of DiffBindFR is not solely due to local ligand energy relaxation and its superior in side chain packing, the performance of other methods is evaluated with stricter energy minimization for the ligand, given the experimental side chain conformation. For poses generated by DiffBindFR, ligand energy minimization is conducted using the side chains as predicted by the model. The energy minimization is performed using the AMBER ff14sb force field 69 for prorein and the Sage force field 70 for ligand in OpenMM 71, as used in the Posebusters paper 47. Fig.4.(A) shows that traditional rigid receptor docking methods like Glide perform best on re-docking when provided with the correct Holo pocket environment, followed by Vina and Gold, with most of their generated docking poses being physically valid. However, their performance significantly deteriorates when docking with Rosetta-repacked proteins, highlighting their heavy dependence on side chain conformations. Traditional flexible docking methods rDock and VinaFlex are also involved in comparison. VinaFlex, heavily reliant on predefined flexible side chains, performs the worst in our scenario where information about flexible side chains is assumed unavailable. rDock, capable of optimizing functional groups prone to forming hydrogen bonds in side chains, achieves higher success rate in repacked proteins compared to Vina and Glide, but lower success rate in proteins with ground-truth side chains. For these traditional methods, their PB-success rate is only slightly lower than their overall success rate, indicating that most generated poses is validated by Posebusters suite due to the physical components in their scoring functions. Therefore, the post ligand optimization using force field does not cause obvious impact to their PB-success rate.
| Method |
PB-success rate |
| Ensemble-CDK2 | Ensemble-EGFR | Ensemble-FXA | ApoRef | CASF2016 | GPCR-AF2 | ||
| Traditional rigid receptor docking methods | Vina | 0.079 | 0.060 | 0.344 | 0.082 | 0.294 | 0.136 |
| LinF9 | 0.061 | 0.119 | 0.365 | 0.089 | 0.272 | 0.076 | |
| Smina | 0.079 | 0.090 | 0.362 | 0.077 | 0.304 | 0.152 | |
| Gnina | 0.099 | 0.090 | 0.394 | 0.082 | 0.320 | 0.106 | |
| Glide | 0.154 | 0.090 | 0.271 | 0.091 | 0.219 | 0.182 | |
| Traditional flexible docking methods | VinaFlex | 0.013 | 0.000 | 0.005 | 0.015 | 0.023 | 0.045 |
| rDock | 0.257 | 0.134 | 0.440 | 0.157 | 0.296 | 0.212 | |
| Deep learning-based docking methods | TankBind-pocket | 0.100 | 0.015 | 0.110 | 0.040 | 0.123 | 0.015 |
| EDM-Dock | 0.051 | 0.015 | 0.009 | 0.011 | 0.064 | 0.00 | |
| KarmaDock Align | 0.135 | 0.045 | 0.009 | 0.047 | 0.136 | 0.015 | |
| DiffBindFR-Smina | 0.564 | 0.403 | 0.789 | 0.434 | 0.566 | 0.227 | |
| DiffBindFR-MDN | 0.674 | 0.478 | 0.794 | 0.476 | 0.636 | 0.212 | |
Best performance in bold for the highest PB-success rate.
Among blind docking methods, DiffDock shows better performance (success rate of 38) than TankBind (16) and EquiBind (2), but most of their generated poses are invalid due to ignoring protein side chains (Fig.4.(A)). Ligand energy minimization significantly improves the PB-success rate of DiffDock (35). TankBind and EquiBind also see improvements in PB-success rate with energy minimization, but still lag behind DiffDock (Fig.4.(B)). Although blind docking is a tough task for its broad searching space in the whole protein, flexible pocket docking method like DiffBindFR, denoising a chaotic side chain conformation into a well-packed conformation having valid interaction with the ligand, has much more objectives for prediction. DiffBindFR, utilizing Smina scoring function or MDN network as the confidence model to select the top-1 pose from 40 generated ones, outperforms all other deep learning-based blind docking methods and pocket docking methods. DiffBindFR-Smina and DiffBindFR-MDN demonstrate both high success rate (50.2 for DiffBindFR-Smina and 48.1 for DiffBindFR-MDN) and PB-success rate (49.1 for DiffBindFR-Smina and 44.4 for DiffBindFR-MDN), with lower penalties by Posebusters compared to other deep learning-based methods, showcasing the capability of DiffBindFR in generating accurate and physically plausible complex poses. The performance of DiffBindFR is comparable to traditional rigid receptor docking methods using known ground-truth side chain conformations for redocking. As is depicted from Supplemental Fig. S7, DiffBindFR shows its effectiveness in binding site identification and pocket side chain recovery on Posebusters test set, as well. Force field optimization has minimal impact on DiffBindFR generated structures, which also demonstrates the high physical plausibility of DiffBindFR generated poses. Among other deep learning-based pocket docking methods, KarmaDock Align achieves the highest success rate (30.4) but a very low PB-success rate (6.1). Force field optimization of ligands rescues most poses with L-RMSD 2 Å into good physical validity. KarmaDock Align, EDM-Dock, TankBind-pocket, Uni-Mol, and EDM-Dock, which focus on fitting the RMSD of the ligand during training, tend to ignore the intra energy of the generated poses and protein side chains, as is shown from supplemental Fig. S9. Indeed, force field optimization is not allowed in realistic docking to meet the demands of high-throughput screening.
Four specific cases from the Posebusters test set (supplementary Fig. S10), never trained or seen by DiffBindFR, are presented to highlight its superiority over other methods focusing solely on ligand coordinates while neglecting ligand conformation validity. In complexes with PDB ID 6TW5, 7PRM, 7T1D, and 7CD9, DiffBindFR successfully docks ligands into precise positions with valid conformations and recovers pocket side chains into good interaction with ligands. In contrast, KarmaDock Align, EDM-Dock, and TankBind-pocket fail to predict correct binding ligand poses, and their generated poses cannot pass the physical plausibility check of the Posebusters suite. As is shown in Supplementary Table. S4, ligand poses generated by EDM-Dock and TankBind-pocket exhibit both internal invalidity (including internal steric clash, bump aromatic ring, etc) and steric clash with proteins, while KarmaDock Align, due to using RDKit for ligand pose alignment, frequently fails in reducing ligand-protein clash.